# Phase Transitions in Anisotropic Turbulence

### v1

@@ I Introduction
Werner Heisenberg, whose foundational contributions to quantum mechanics were recognized with the Nobel Prize in Physics, is said to have uttered on his deathbed that, if he were allowed to ask God two questions, they would be “Why relativity?" and "Why turbulence?” and he was allegedly sure that God would be able to answer the former question. Turbulence here refers to a state of fluid flows realized when the Reynolds number $\rm Re$ (the ratio of the magnitude of inertial to viscous forces) attains large values, leading to strongly nonlinear, multi-scale, chaotic dynamics. Although the quote may be apocryphal, Heisenberg did indeed devote significant efforts to the theoretical study of turbulenceHeisenberg (1985) in three spatial dimensions. However, his efforts were not crowned with success – it was Andrey Kolmogorov who, in 1941, proposed the first largely successful attempt at a quantitative theoryKolmogorov (1941) of the forward energy cascade transferring kinetic energy from large forcing scales to small, dissipative scales in three-dimensional (3D) turbulence, now understood to be a consequence of nonlinear vortex stretchingFrisch (1995) and strain self-amplification.Johnson (2020, 2021)
@@
In the 1950s, there were significant advances in numerical weather prediction,Charney, Fjørtoft, and von Neumann (1950) realizing a vision formulated decades earlier by Lewis Fry Richardson.Richardson (1922) This progress relied on the first available digital computers to solve the barotropic, i.e. two-dimensional (2D), vorticity equation, an idealized model of large-scale motions in the Earth’s atmosphere. Around that time, there were first inklings Fjørtoft (1953) that the properties of such turbulence restricted to two spatial dimensions were fundamentally different from its 3D counterpart due to the simultaneous conservation of kinetic energy and enstrophy (mean squared vorticity). The full implications of this observation were first laid out in the seminal contributions of Robert Kraichnan Kraichnan (1967, 1971) (complemented by the works of LeithLeith (1968) and BatchelorBatchelor (1969)), establishing that turbulence in 2D fluids exhibits a dual cascade, with kinetic energy being transferred towards larger scales (an inverse energy cascade), and enstrophy being transferred towards smaller scales (a forward enstrophy cascade) from a given stirring scale. Fluid flow in two dimensions is an adequate (albeit often approximate) description for many systems where one spatial dimension is strongly constrained, such as in a thin-layer geometry or in the presence of some other physical mechanism, like rotation or strong magnetic fields, preventing variations in the velocity field along one dimension. This situation is in fact encountered across physical scales from atomic through planetary. At the smallest scales, strongly correlated electron systems in ultra-pure materials such as graphene exhibit 2D fluid-like behavior.Levitov and Falkovich (2016); Bandurin et al. (2018); Keser et al. (2021); Narozhny (2022) At similarly small scales, another striking realization of 2D flow is found in paraxial fluids of light.Ferreira et al. (2022); Baker-Rasooli et al. (2023); Congy et al. (2024) Yet another example of 2D flows is met in Bose-Einstein condensates.Horng et al. (2009); Reeves, Anderson, and Bradley (2012); Seo et al. (2017); Gauthier et al. (2019); Johnstone et al. (2019) At larger scales, it has been shown that a 2D equation of motion provides a good approximation for flows in thin films of dense bacterial suspensions Wu and Libchaber (2000); Sokolov et al. (2007); Kurtuldu et al. (2011); Wensink et al. (2012) as well as flows in soap films.Martin et al. (1998); Rivera, Vorobieff, and Ecke (1998); Vorobieff, Rivera, and Ecke (1999); Kellay and Goldburg (2002); Shakeel and Vorobieff (2007) At yet larger scales, flows in plasmas subject to strong magnetic fields, like those occurring in Tokamak devices, also display a near-2D structure.Xia and Shats (2003); Fujisawa (2008) Even at the large scales of planetary atmospheres, flows are typically constrained to be approximately 2D due to the combined effects of geometric confinement, planetary rotation and density stratification.Byrne and Zhang (2013); King, Vogelzang, and Stoffelen (2015); Young and Read (2017); Siegelman et al. (2022); Alexakis et al. (2024) The literature on 2D turbulence is extensive, and this mini-review is not intended to provide an exhaustive description. Instead, we point the reader to existing review articles on the topic.Kraichnan and Montgomery (1980); Tabeling (2002); Kellay and Goldburg (2002); Boffetta and Ecke (2012)
@@
A noteworthy feature of 2D turbulence is that in a finite system the inverse cascade can cause energy to pile up at the largest scales, thereby producing large-scale Bose condensation, first observed in experiments by SommeriaSommeria (1986) and later in simulations by Smith and YakhotSmith and Yakhot (1993) (see also the theoretical work by Falkovich Falkovich (1992)), which is associated with the emergence of coherent flow structures such as large-scale vortices or unidirectional jets. The possibility of such condensation had already been suggested earlier by Kraichnan,Kraichnan (1967) who proposed the idea that equilibrium statistical mechanics based on energy and enstrophy conservation might be used to describe the condensate, building on related earlier work by Lee Lee (1952) and Onsager.Onsager (1949) The later groundbreaking works by Sommeria Robert and Sommeria (1991) and MillerMiller, Weichman, and Cross (1992) in the 1990s significantly refined these approaches, leading to detailed statistical mechanical descriptions of the large-scale flows based on the 2D Euler equation of ideal (i.e., inviscid and unforced) fluids. This sparked a number studies on the statistical mechanics of two-dimensional and geophysical flows, reviewed by Bouchet and Venaille.Bouchet and Venaille (2012) Complementing the statistical mechanics approach which relies on the Euler equations of ideal fluids, a significant body of recent work has established a detailed characterization of forced-dissipative large-scale condensates within 2D turbulence using a combination of direct numerical simulations (DNS) and theoretical approaches.Chertkov et al. (2007); Laurie et al. (2014); Frishman (2017); Frishman and Herbert (2018); Doludenko et al. (2021)
@@
While 2D turbulence has attracted great interest in its own right, in most cases it remains only an approximation for real physical systems. Even when close to 2D, most realistic fluid flows (such as the Earth’s atmospheric circulation which is largely confined within the weather layer whose depth of $\sim 10{\rm km}$ is small compared to the horizontal size of storms of $\sim 1000{\rm km}$) generically retain some degree of three-dimensionality. Specifically in the atmosphere, an important source of such 3D motions is (moist) convection. Their presence motivates the study of the broader class of quasi-2D flows, where the fluid motion is constrained to occur primarily in two dimensions, but where variations along the third dimension (referred to as the vertical in the following) may be present. These vertical variations can strongly impact the dynamics, since they introduce vortex stretching which breaks exact enstrophy conservation, and thus the nature of the resulting turbulent cascades is non-trivial. While quasi-2D turbulence is a problem of long-standing interest, studied in early simulations since the late 1990s Smith, Chasnov, and Waleffe (1996); Celani, Musacchio, and Vincenzi (2010) (see also the review in Ref. Danilov and Gurarie (2000)), the last few years in particular have seen substantial progress in our understanding of this important class of fluid flows. In particular, it has been realized that quasi-2D turbulence features a bidirectional (or split) energy cascade, where energy injected at some intermediate forcing scale simultaneously cascades both to larger and to smaller scales at rates depending on physical control parameters, e.g. the depth of a thin fluid layer, the planetary rotation rate or the strength of density stratification. A bidirectional cascade in an anisotropic flow is facilitated by the multi-scale nature of turbulence: large-scales can be strongly constrained to exhibit approximately 2D dynamics, while small scales are unrestrained, evolving as fully 3D.
@@
We emphasize that this review is focused on transitions between distinct turbulent flow states, with energy transfers across a wide range of scales (implying that the system is out of statistical equilibrium) on both sides of the transition. This problem is qualitatively distinct from the transition between laminar flow (which does not feature energy fluxes across scales) and turbulence. This mini-review reflects a subjective choice of topics that the author deems of the greatest interest. Alternative and complementary perspectives are presented in other recently published reviews.Alexakis and Biferale (2018); Alexakis (2023); Boffetta (2023); Marston and Tobias (2023)
The remainder of this paper is structured as follows. In Section II, we give background information on the basic properties of 2D and 3D turbulence. In Section III, we summarize recent findings on ‘dimensional’ transitions between 3D, quasi-2D and 2D turbulence in different physical systems, with an emphasis on geo- and astrophysically relevant effects including thin-layer geometry, rotation and density stratification. In Section IV, we describe examples of anisotropic flows displaying multistability and transitions between distinct turbulent flow states induced by turbulent fluctuations. Next, in Section V, we discuss phase transitions recently identified within variants of 2D turbulence and finally, in Section VI, we conclude by putting these developments in a broader context.
@@ II.1 Three-dimensional turbulence
where $P$ is the pressure divided by the constant density, $\nu$ is the kinematic viscosity and $\mathbf{f}$ is an externally imposed body force injecting power $\epsilon$ on a well-defined forcing scale $\ell_{f}=2\pi/k_{f}$ (analogous to the size of a stirring device), with the corresponding forcing wave number $k_{f}$. The term $-\alpha\overline{\mathbf{u}}$, where $\overline{(\cdot)}$ denotes the vertical average, is an additional “bottom drag” modeling the effects of boundaries. We consider a Cartesian domain of dimensions $L\times L\times H$, subject to periodic boundary conditions for simplicity where $H$ (assumed to be less than $L$) is varied to study the suppression of 3D variations in the flow.
@@
$E(k)=\frac{1}{2\Delta k}\sum_{k\leq\mathbf{p}\leq k+\Delta k}|\hat{\mathbf{u}}_{\mathbf{k}}|^{2},$ (4)
where $\Delta k=2\pi/L$ is the minimum wave number present in the system. The spectrum is related to the total energy by $\mathcal{E}=\sum_{k}E(k)\Delta k$, with $E(k)\Delta k$ being the energy contained in the wave number interval $[k,k+\Delta k]$.
@@
up to a cut-off wave number (now named after Kolmogorov) where viscous dissipation terminates the forward energy cascade, given by $k_{\nu}=\epsilon^{1/4}/\nu^{3/4}=k_{f}{\rm Re}^{3/4}$, and $C_{K}$ is the Kolmogorov constant. At wave numbers smaller than $k_{f}$ (corresponding to scales larger than $\ell_{f}$), there is no flux of energy, and thus these scales are expected to reach a thermal equilibrium state characterized by an equipartition of energy among modes.Dallas, Fauve, and Alexakis (2015); Alexakis and Brachet (2019); Gorce and Falcon (2022) Later work revealed that the forward energy cascade is not local in scale, and the probability distribution of velocity differences between different points is non-Gaussian, developing heavy tails at small scales. This phenomenon, which is known as intermittency, is associated with a small deviation from the $-5/3$ spectral exponent and has received much attention, but will not be discussed further here. A key physical mechanism behind the forward energy cascade of 3D turbulence is vortex stretching,Tennekes and Lumley (1972) i.e. vorticity amplification at smaller scales by along-vortex velocity gradients, which is associated with the term $\bm{\omega}\cdot\nabla\mathbf{u}$ in the evolution equation for vorticity $\bm{\omega}$, in addition to strain amplification.Johnson (2020, 2021) Vortex stretching leads to a small-scale filamentary structure which can be seen in the vorticity iso-contours from a high-resolution 3D homogeneous and isotropic turbulence simulation, shown in Fig. 1.
Figure 1: Iso-contours of vorticity from a high-resolution simulation of 3D homogeneous and isotropic turbulence. The elongated small-scale vorticity filaments visible in the snapshot result from vortex stretching. Figure adapted from McKeown et al. 2023.(McKeown et al., 2023)
@@ II.2 Two-dimensional turbulence
where $f_{\omega_{z}}=\bm{\hat{z}}\cdot(\nabla\times\mathbf{f})$. Note that there is no vortex stretching term on the right-hand side of Eq. (6) since the vorticity and velocity vectors are perpendicular for 2D flow. This means that a key mechanism responsible for the forward energy cascade is absent, with profound implications for energy transfers. Equation (6) has two ideal quadratic invariants: energy $E=\langle\frac{1}{2}|\mathbf{u}|^{2}\rangle$ and enstrophy $\Omega=\langle\omega_{z}^{2}\rangle$. Moreover, the Casimir invariants $\langle g(\omega_{z})\rangle$ are conserved for any differentiable function $g$, which presents a challenge in particular for statistical mechanics approaches.Bouchet and Venaille (2012) In a statistically steady state, dissipation balances the injection of energy and enstrophy i.e.
@@
where $\epsilon$ is again the injected power, associated with an enstrophy injection rate $\eta$, the energy dissipation rates are given by $D^{\mathcal{E}}_{\alpha}=\alpha\langle\mathbf{u}^{2}\rangle$, $D^{\mathcal{E}}_{\nu}=\nu\langle\omega_{z}^{2}\rangle$ and the enstrophy dissipation rates by $D^{\Omega}_{\alpha}=\alpha\langle\omega_{z}^{2}\rangle$ and $D_{\nu}^{\Omega}=\nu\langle|\nabla\omega_{z}|^{2}\rangle$.
Following his involvementCharney, Fjørtoft, and von Neumann (1950) in early numerical weather forecasting based on 2D models, FjørthoftFjørtoft (1953) proposed an argument indicating that kinetic energy cannot be transferred from large to small scales in 2D and must instead be transferred inversely, i.e. from small to large scales. His argument is based on the observation that the enstrophy spectrum $\Omega(k)$ is directly related to the energy spectrum by $\Omega(k)=2k^{2}E(k)$. Fjørthoft’s argument considers an ideal fluid and assumes that energy is initially concentrated in some wave number interval of width $\delta k=\left(\mathcal{E}^{-1}\sum_{k}(k-\overline{k}_{\mathcal{E}})^{2}E(k)\Delta k\right)^{1/2}$, where $\overline{k}_{\mathcal{E}}=\mathcal{E}^{-1}\sum_{k}kE(k)\Delta k$ is the energy centroid. It can easily be shown that $(\delta k)^{2}=\Omega/(2\mathcal{E})-\overline{k}_{\mathcal{E}}^{2}$. Since $\Omega/(2\mathcal{E})$ is conserved and nonlinear interactions tend to increase $\delta k$, the energy centroid $\overline{k}_{\mathcal{E}}$ must decrease, i.e. energy is transferred to larger scales. Similarly defining the enstrophy centroid as $\overline{k}_{\Omega}=\Omega^{-1}\sum_{k}k\Omega(k)\Delta k$, it follows (using the Cauchy-Schwarz inequality) that $\overline{k}_{\mathcal{E}}\overline{k}_{\Omega}\geq\Omega/(2\mathcal{E})$. This implies that as $\overline{k}_{\mathcal{E}}$ decreases, $\overline{k}_{\Omega}$ necessarily increases, indicating a forward transfer of enstrophy to smaller scales. The resulting balance relations for a dual inverse energy forward enstrophy cascade, in an infinite domain and in the limit of large $\rm Re$, $\rm Re_{\alpha}$, read
@@
This dual cascade picture deduced from Fjørthoft’s argument, and later fully developed by Kraichnan,Kraichnan (1967, 1971) has been corroborated in high-resolution numerical simulations.Boffetta (2007); Boffetta and Musacchio (2010) By similar arguments as in the 3D case, one finds that the associated energy spectrum at $k_{\alpha}\ll k\ll k_{f}$ (in the inverse energy cascade range) is of the form $E(k)\propto k^{-5/3}$, while at $k_{f}\ll k\ll k_{\nu}$ (in the forward enstrophy cascade range) it is steeper $E(k)\propto k^{-3}$. The wave number of frictional arrest of the inverse energy cascade is given by $k_{\alpha}=\alpha^{3/2}/\epsilon^{1/2}=k_{f}{\rm Re_{\alpha}}^{-3/2}$, while the wave number of viscous enstrophy dissipation reads $k_{\nu}=\eta^{1/6}/\nu^{1/2}=k_{f}{\rm Re}^{1/2}$. Note that the $\rm Re$-dependence of $k_{\nu}$ is different in 2D and 3D.

### v2

@@ I Introduction
Werner Heisenberg, whose foundational contributions to quantum mechanics were recognized with the Nobel Prize in Physics, is said to have uttered on his deathbed that, if he were allowed to ask God two questions, they would be “Why relativity?” and “Why turbulence?” and he was allegedly sure that God would be able to answer the former question. Turbulence here refers to a state of fluid flows realized when the Reynolds number $\rm Re$ (the ratio of the magnitudes of inertial to viscous forces) attains large values, leading to strongly nonlinear, multi-scale, chaotic dynamics. Although the quote may be apocryphal, Heisenberg did indeed devote significant efforts to the theoretical study of turbulenceHeisenberg (1985) in three spatial dimensions. However, his efforts were not crowned with success – it was Andrey Kolmogorov who, in 1941, proposed the first largely successful attempt at a quantitative theoryKolmogorov (1941) of the forward energy cascade transferring kinetic energy from large forcing scales to small, dissipative scales in three-dimensional (3D) turbulence, now understood to be a consequence of nonlinear vortex stretchingFrisch (1995) and strain self-amplification.Johnson (2020, 2021)
@@
In the 1950s, there were significant advances in numerical weather prediction,Charney, Fjørtoft, and von Neumann (1950) realizing a vision formulated decades earlier by Lewis Fry Richardson.Richardson (1922) This progress relied on the first available digital computers to solve the barotropic, i.e. two-dimensional (2D), vorticity equation, an idealized model of large-scale motions in the Earth’s atmosphere. Around that time, there were first inklings Fjørtoft (1953) that the properties of such turbulence restricted to two spatial dimensions were fundamentally different from its 3D counterpart due to the simultaneous conservation of kinetic energy and enstrophy (mean squared vorticity). The full implications of this observation were first laid out in the seminal contributions of Robert Kraichnan Kraichnan (1967, 1971) (complemented by the works of LeithLeith (1968) and BatchelorBatchelor (1969)), establishing that turbulence in 2D fluids exhibits a dual cascade, with kinetic energy being transferred towards larger scales (an inverse energy cascade), and enstrophy being transferred towards smaller scales (a forward enstrophy cascade) from a given stirring scale. Fluid flow in two dimensions is an adequate (albeit often approximate) description for many systems where one spatial dimension is strongly constrained, such as in a thin-layer geometry or in the presence of some other physical mechanism, like rotation or strong magnetic fields, preventing variations in the velocity field along one dimension. This situation is in fact encountered across physical scales from atomic through planetary. At the smallest scales, strongly correlated electron systems in ultra-pure materials such as graphene exhibit 2D fluid-like behaviorLevitov and Falkovich (2016); Bandurin et al. (2018); Keser et al. (2021); Narozhny (2022) and while the flows in this context are typically highly viscous and non-turbulent, vortices have been observed in such electron fluids.Aharon-Steinberg et al. (2022) At similarly small scales, another striking realization of 2D flow is found in paraxial fluids of light.Ferreira et al. (2022); Baker-Rasooli et al. (2023); Congy et al. (2024) Yet another example of 2D flows is met in Bose-Einstein condensates.Horng et al. (2009); Reeves, Anderson, and Bradley (2012); Seo et al. (2017); Gauthier et al. (2019); Johnstone et al. (2019) At larger scales, it has been shown that 2D equations of motion provide a good approximation for flows in thin films of dense bacterial suspensions Wu and Libchaber (2000); Sokolov et al. (2007); Kurtuldu et al. (2011); Wensink et al. (2012) as well as flows in soap films.Martin et al. (1998); Rivera, Vorobieff, and Ecke (1998); Vorobieff, Rivera, and Ecke (1999); Kellay and Goldburg (2002); Shakeel and Vorobieff (2007) At yet larger scales, flows in plasmas subject to strong magnetic fields, like those occurring in Tokamak devices, also display a near-2D structure.Xia and Shats (2003); Fujisawa (2008) Even at the large scales of planetary atmospheres, flows are typically constrained to be approximately 2D due to the combined effects of geometric confinement, planetary rotation and density stratification.Byrne and Zhang (2013); King, Vogelzang, and Stoffelen (2015); Young and Read (2017); Siegelman et al. (2022); Alexakis et al. (2024) The literature on 2D turbulence is extensive, and this mini-review is not intended to provide an exhaustive description. Instead, we point the reader to existing review articles on the topic.Kraichnan and Montgomery (1980); Tabeling (2002); Kellay and Goldburg (2002); Boffetta and Ecke (2012)
@@
A noteworthy feature of 2D turbulence is that in a finite system the inverse cascade can cause energy to pile up at the largest scales, thereby producing large-scale Bose condensation, first observed in experiments by SommeriaSommeria (1986) and later in simulations by Smith and YakhotSmith and Yakhot (1993) (see also the theoretical work by Falkovich Falkovich (1992)), which is associated with the emergence of coherent flow structures such as large-scale vortices or unidirectional jets. The possibility of such condensation had already been suggested earlier by Kraichnan,Kraichnan (1967) who proposed the idea that equilibrium statistical mechanics based on energy and enstrophy conservation might be used to describe the condensate, building on related earlier work by Lee Lee (1952) and Onsager.Onsager (1949) The later groundbreaking works by Sommeria Robert and Sommeria (1991) and MillerMiller, Weichman, and Cross (1992) in the 1990s significantly refined these approaches, leading to detailed statistical mechanical descriptions of the large-scale flows based on the 2D Euler equation of ideal (i.e., inviscid and unforced) fluids. This sparked a number of studies on the statistical mechanics of two-dimensional and geophysical flows, reviewed by Bouchet and Venaille.Bouchet and Venaille (2012) Complementing the statistical mechanics approach which relies on the Euler equations of ideal fluids, a significant body of recent work has established a detailed characterization of forced-dissipative large-scale condensates within 2D turbulence using a combination of direct numerical simulations (DNS) and theoretical approaches.Chertkov et al. (2007); Laurie et al. (2014); Frishman (2017); Frishman and Herbert (2018); Doludenko et al. (2021) We stress that the term 2D flow is used here to refer specifically to planar flow (with two components), which should distinguished from two-dimensional three-component (2D3C) flow, sometimes also referred to as 2.5-dimensional flow, where the velocity field only depends on two coordinates, but retains three components. For instance, even when the flow becomes vertically invariant, e.g., in a very thin layer or under the influence of rapid rotation, the vertical component of velocity obeys a passive scalar equation with an associated forward energy cascade (if the domain has periodic boundary conditions in the vertical). A problem for which this distinction is important is the magnetohydrodynamic dynamo instability: 2D3C flows can induce magnetic fields, i.e. dynamo action,Seshasayanan and Alexakis (2016); Seshasayanan, Gallet, and Alexakis (2017) while planar flows cannot.ZELDOVICH and Ruzmaikin (1980) The transition between 2D flow with two components and 2D3C flow has recently been investigated in a toroidal geometryAgoua et al. (2024), but will not be discussed further here.
@@
While 2D turbulence has attracted great interest in its own right, in most cases it remains only an approximation for real physical systems. Even when close to 2D, most realistic fluid flows (such as the Earth’s atmospheric circulation which is largely confined within the troposphere whose depth of $\sim 10{\rm km}$ is small compared to the horizontal size of storms of $\sim 1000{\rm km}$) generically retain some degree of three-dimensionality. Specifically in the atmosphere, an important source of such 3D motions is (moist) convection. Their presence motivates the study of the broader class of quasi-2D flows, where the fluid motion is constrained to occur primarily in two dimensions, but where variations along the third dimension (referred to as the vertical in the following) may be present. These vertical variations can strongly impact the dynamics, since they introduce vortex stretching which breaks exact enstrophy conservation, and thus the nature of the resulting turbulent cascades is non-trivial. While quasi-2D turbulence is a problem of long-standing interest, studied in early simulations since the late 1990s Smith, Chasnov, and Waleffe (1996); Celani, Musacchio, and Vincenzi (2010) (see also the review in Ref. Danilov and Gurarie (2000)), the last few years in particular have seen substantial progress in our understanding of this important class of fluid flows. In particular, it has been realized that quasi-2D turbulence features a bidirectional (or split) energy cascade, where energy injected at some intermediate forcing scale simultaneously cascades both to larger and to smaller scales at rates depending on physical control parameters, e.g. the depth of a thin fluid layer, the planetary rotation rate or the strength of density stratification. A bidirectional cascade in an anisotropic flow is facilitated by the multi-scale nature of turbulence: large-scales can be strongly constrained to exhibit approximately 2D dynamics, while small scales are unrestrained, evolving as fully 3D. The upscale transfer of kinetic energy is observed in tropical cyclone atmospheric boundary layer dynamicsShao, Zhang, and Tang (2023), where it leads to reduced drag, which can lead to forecast errors if not parameterized correctly.
@@
We emphasize that this review is focused on transitions between distinct turbulent flow states, with energy transfers across a wide range of scales (implying that the system is out of statistical equilibrium) on both sides of the transition. This problem is qualitatively distinct from the transition between laminar flow (which does not feature energy fluxes across scales) and turbulence. This mini-review reflects a subjective choice of topics that the author deems of the greatest interest. Alternative and complementary perspectives are presented in other recently published reviews.Alexakis and Biferale (2018); Alexakis (2023); Boffetta (2023); Marston and Tobias (2023) The remainder of this paper is structured as follows. In Section II, we give background information on the basic properties of 2D and 3D turbulence. In Section III, we summarize recent findings on ‘dimensional’ transitions between 3D, quasi-2D and 2D turbulence in different physical systems, with an emphasis on geo- and astrophysically relevant effects including thin-layer geometry, rotation and density stratification. In Section IV, we describe examples of anisotropic flows displaying multistability and transitions between distinct turbulent flow states induced by turbulent fluctuations. Next, in Section V, we discuss phase transitions recently identified within variants of 2D turbulence and finally, in Section VI, we conclude by putting these developments in a broader context.
@@ II.1 Three-dimensional turbulence
where $P$ is the pressure divided by the constant density, $\nu$ is the kinematic viscosity and $\mathbf{f}$ is an externally imposed body force injecting power $\epsilon$ on a well-defined forcing scale $\ell_{f}=2\pi/k_{f}$ (analogous to the size of a stirring device), with the corresponding forcing wave number $k_{f}$. The term $-\alpha\overline{\mathbf{u}}$, where $\overline{(\cdot)}$ denotes the vertical average, is an additional “bottom drag” modeling the effects of boundaries. We consider a Cartesian domain of dimensions $L\times L\times H$, subject to triply periodic boundary conditions for simplicity, where $H$ (assumed to be less than $L$) is varied to study the suppression of 3D variations in the flow. All results discussed in the following assume periodicity in the vertical direction, unless otherwise stated. Alternative, e.g., no-slip boundary conditions lead to the formation of boundary layers which may impact the flow, but this is not considered here.
@@
$E(k)=\frac{1}{2\Delta k}\sum_{\stackrel{{\scriptstyle\mathbf{p}}}{{k\leq|\mathbf{p}|\leq k+\Delta k}}}|\hat{\mathbf{u}}_{\mathbf{p}}|^{2},$ (4)
where $\Delta k=2\pi/L$ is the minimum wave number present in the system. The spectrum is related to the total energy by $\mathcal{E}=\sum_{k}E(k)\Delta k$, with $E(k)\Delta k$ being the energy contained in the wave number interval $[k,k+\Delta k]$. Based on the fraction of energy $p_{k}=E(k)\Delta k/\mathcal{E}$ contained in each wave number shell, Verma al.Verma and Chatterjee (2022); Verma, Stepanov, and Delache (2024) recently introduced hydrodynamic entropy to quantify the multiscale disorder of turbulence (not discussed further here).
@@
up to a cut-off wave number (now named after Kolmogorov) where viscous dissipation terminates the forward energy cascade, given by $k_{\nu}=\epsilon^{1/4}/\nu^{3/4}=k_{f}{\rm Re}^{3/4}$, and $C_{K}$ is the Kolmogorov constant. At wave numbers smaller than $k_{f}$ (corresponding to scales larger than $\ell_{f}$), there is no flux of energy, and thus these scales are expected to reach a thermal equilibrium state characterized by an equipartition of energy among modes.Dallas, Fauve, and Alexakis (2015); Alexakis and Brachet (2019); Gorce and Falcon (2022) Later work revealed Frisch (1995) that the forward energy cascade is not local in scale, and the probability distribution of velocity increments between different points in space is non-Gaussian, developing heavy tails at small scales. This phenomenon, which is known as intermittency, is associated with a small deviation from the $-5/3$ spectral exponent and has received much attention, but will not be discussed further here. A key physical mechanism facilitating the forward energy cascade of 3D turbulence is vortex stretching,Tennekes and Lumley (1972) i.e. vorticity amplification at smaller scales by along-vortex velocity gradients, which is associated with the term $\bm{\omega}\cdot\nabla\mathbf{u}$ in the evolution equation for vorticity $\bm{\omega}$, in addition to strain amplification.Johnson (2020, 2021) Vortex stretching leads to a small-scale filamentary structure which can be seen in the vorticity iso-contours from a high-resolution 3D homogeneous and isotropic turbulence simulation, shown in Fig. 1.
Figure 1: Iso-contours of vorticity from a high-resolution simulation of 3D homogeneous and isotropic turbulence. The elongated small-scale vorticity filaments visible in the snapshot result from vortex stretching. Source: Ryan McKeown et al. New J. Phys. 2023, 25 103029; licensed under a Creative Commons Attribution (CC BY) license.
@@ II.2 Two-dimensional turbulence
where $f_{\omega_{z}}=\bm{\hat{z}}\cdot(\nabla\times\mathbf{f})$. Note that there is no vortex stretching term on the right-hand side of Eq. (6) since the vorticity and velocity vectors are perpendicular for planar 2D flow. This means that a key mechanism responsible for the forward energy cascade is absent, with profound implications for energy transfers. Equation (6) has two ideal quadratic invariants: energy $E=\langle\frac{1}{2}|\mathbf{u}|^{2}\rangle$ and enstrophy $\Omega=\langle\omega_{z}^{2}\rangle$. Moreover, the Casimir invariants $\langle g(\omega_{z})\rangle$ are conserved for any differentiable function $g$, which presents a challenge in particular for statistical mechanics approaches.Bouchet and Venaille (2012) In a statistically steady state, dissipation balances the injection of energy and enstrophy i.e.
@@
where $\epsilon$ is again power injected at the forcing scale, associated with an enstrophy injection rate $\eta$, the energy dissipation rates are given by $D^{\mathcal{E}}_{\alpha}=\alpha\langle\mathbf{u}^{2}\rangle$, $D^{\mathcal{E}}_{\nu}=\nu\langle\omega_{z}^{2}\rangle$ and the enstrophy dissipation rates by $D^{\Omega}_{\alpha}=\alpha\langle\omega_{z}^{2}\rangle$ and $D_{\nu}^{\Omega}=\nu\langle|\nabla\omega_{z}|^{2}\rangle$.
Following his involvementCharney, Fjørtoft, and von Neumann (1950) in early numerical weather forecasting based on 2D models, FjørtoftFjørtoft (1953) proposed an argument indicating that kinetic energy cannot be transferred from large to small scales in 2D and must instead be transferred inversely, i.e. from small to large scales. His argument is based on the observation that the enstrophy spectrum $\Omega(k)$ is directly related to the energy spectrum by $\Omega(k)=2k^{2}E(k)$. Fjørtoft’s argument considers an ideal fluid and assumes that energy is initially concentrated in some wave number interval of width $\delta k=\left(\mathcal{E}^{-1}\sum_{k}(k-\overline{k}_{\mathcal{E}})^{2}E(k)\Delta k\right)^{1/2}$, where $\overline{k}_{\mathcal{E}}=\mathcal{E}^{-1}\sum_{k}kE(k)\Delta k$ is the energy centroid. It can easily be shown that $(\delta k)^{2}=\Omega/(2\mathcal{E})-\overline{k}_{\mathcal{E}}^{2}$. Since $\Omega/(2\mathcal{E})$ is conserved and nonlinear interactions tend to increase $\delta k$, the energy centroid $\overline{k}_{\mathcal{E}}$ must decrease, i.e. energy is transferred to larger scales. Similarly defining the enstrophy centroid as $\overline{k}_{\Omega}=\Omega^{-1}\sum_{k}k\Omega(k)\Delta k$, it follows (using the Cauchy-Schwarz inequality) that $\overline{k}_{\mathcal{E}}\overline{k}_{\Omega}\geq\Omega/(2\mathcal{E})$. This implies that as $\overline{k}_{\mathcal{E}}$ decreases, $\overline{k}_{\Omega}$ necessarily increases, indicating a forward transfer of enstrophy to smaller scales. The resulting balance relations for a dual inverse energy forward enstrophy cascade, in an infinite domain and in the limit of large $\rm Re$, $\rm Re_{\alpha}$, read
@@
This dual cascade picture deduced from Fjørtoft’s argument, and later fully developed by Kraichnan,Kraichnan (1967, 1971) has been corroborated in high-resolution numerical simulations.Boffetta (2007); Boffetta and Musacchio (2010) By similar arguments as in the 3D case, one finds that in 2D the associated energy spectrum at $k_{\alpha}\ll k\ll k_{f}$ (in the inverse energy cascade range) is of the form $E(k)\propto k^{-5/3}$, while at $k_{f}\ll k\ll k_{\nu}$ (in the forward enstrophy cascade range) it is steeper $E(k)\propto k^{-3}$. The wave number of frictional arrest of the inverse energy cascade is given by $k_{\alpha}=\alpha^{3/2}/\epsilon^{1/2}=k_{f}{\rm Re_{\alpha}}^{-3/2}$, while the wave number of viscous enstrophy dissipation reads $k_{\nu}=\eta^{1/6}/\nu^{1/2}=k_{f}{\rm Re}^{1/2}$. Note that the $\rm Re$-dependence of $k_{\nu}$ is different in 2D and 3D.